# Emulation Capacity between Idempotent Channels

### v1

@@ Acknowledgements
We thank Robert Salzmann for helpful discussions about [9] and [10], as well as about sets of approximate fixed point of channels. ID and OF acknowledge funding by the European Research Council (ERC Grant AlgoQIP, Agreement No. 851716). LG acknowledges funding by National Natural Science Foundation of China (NNFSC) via grant No.12401163. This work started when MR was a Marie Skłodowska-Curie Fellow at ENS de Lyon, and he acknowledges funding from the European Union’s Horizon Research and Innovation programme, grant Agreement No. HORIZON-MSCA-2022-PF-01 (Project number: 101108117).
@@ A.2 Approximate Multiplicative Domains
As $V=v+v^{*}$ is Hermitian, we have $V=UDU^{*}$ with $U$ a unitary matrix and $D$ a diagonal matrix with real coefficients. Conjugating the last inequality by $U$, we get:
@@ A.3 Proof of Theorem 10
$\|\mathcal{F}\|_{\text{cb}}=\sup_{n\in\mathbb{N}^{*}}\|\mathcal{F}\otimes\mathrm{Id}_{n}\|_{\infty\rightarrow\infty},$
@@
$\|\mathcal{F}\|_{\diamond}=\|\mathcal{F}^{*}\|_{\text{cb}}.$
@@
$\delta\geq\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}=\|\mathcal{F}^{*}-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}\|_{\text{cb}}.$
@@
$\leq\|\mathcal{F}^{*}-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}\|_{cb}\|x\|$
@@
$\|\Phi(x^{*}x)-x^{*}x\|=\|\Phi^{*}(x^{*}x)-\mathcal{F}^{*}(x^{*}x)\|\leq\|\Phi^{*}-\mathcal{F}^{*}\|_{\text{cb}}\|x^{*}x\|\leq\delta\|x\|^{2}.$
@@
Let $\delta\geq\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}.$ By Lemmas 1415 and 17, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is a $\max\{\delta,\frac{6d\delta}{\lambda_{\min,1}}\}$-inclusion (we trivially have that $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is Hermitian preserving). However, as
@@
Basically, we simply apply Proposition 7 to $\mathcal{F}^{\otimes k}:\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes k})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes k})$} and $\mathcal{G}^{\otimes n}:\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})$. Just note that the dimension of $\mathcal{H}_{\mathcal{F}}^{\otimes k}$ is $d^{k}$, with $d=\dim(\mathcal{H}_{\mathcal{F}})$. Thus, by Proposition 7, if the hypotheses of Theorem 10 are met, for all $p\in[1,+\infty]$, we have:
$\|\lambda(\mathcal{F}^{\otimes k})\|_{p}\leq\|\lambda(\mathcal{G}^{\otimes n})\|_{p}.$
This imply, by the multiplicativity of the $\ell_{p}$-norms of the shape vector with the tensor product, that

### v2

@@ Acknowledgements
We thank Robert Salzmann for helpful discussions about [10] and [11], as well as about sets of approximate fixed points of channels. We thank Satvik Singh and Arturo Konderak for pointing out a mistake in Proposition 1 in the first version of the present article. We also thank Arturo Konderak for drawing our attention to [1]. ID and OF acknowledge funding by the European Research Council (ERC Grant AlgoQIP, Agreement No. 851716). LG acknowledges funding by National Natural Science Foundation of China (NNFSC) via grant No.12401163. This work started when MR was a Marie Skłodowska-Curie Fellow at ENS de Lyon, and he acknowledges funding from the European Union’s Horizon Research and Innovation programme, grant Agreement No. HORIZON-MSCA-2022-PF-01 (Project number: 101108117).
@@ A.2 Approximate Multiplicative Domains
As $V=v+v^{*}$ is Hermitian, we have $V=UDU^{*}$ with $U$ a unitary matrix and $D$ a diagonal matrix with real coefficients. Conjugating the last inequality by $U$, we get
@@ A.3 Proof of Theorem 11
$\|\mathcal{F}\|_{\mathrm{cb}}=\sup_{n\in\mathbb{N}^{*}}\|\mathcal{F}\otimes\mathrm{Id}_{n}\|_{\infty\rightarrow\infty},$
@@
$\|\mathcal{F}\|_{\diamond}=\|\mathcal{F}^{*}\|_{\mathrm{cb}}.$
@@
$\delta\geq\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}=\|\mathcal{F}^{*}-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}\|_{\mathrm{cb}}.$
@@
$\leq\|\mathcal{F}^{*}-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}\|_{\mathrm{cb}}\|x\|$
@@
$\|\Phi(x^{*}x)-x^{*}x\|=\|\Phi^{*}(x^{*}x)-\mathcal{F}^{*}(x^{*}x)\|\leq\|\Phi^{*}-\mathcal{F}^{*}\|_{\mathrm{cb}}\|x^{*}x\|\leq\delta\|x\|^{2}.$
@@
Let $\delta\geq\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}$. Since we assumed that $\mathcal{F}$ and $\mathcal{G}$ are equal to their reduced channels, i.e. $\mathcal{F}=\widehat{\mathcal{F}}$ and $\mathcal{G}=\widehat{\mathcal{G}}$, $\operatorname{Rg}(\mathcal{F}^{*})$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ and $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$. By Lemmas 1516 and 18, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is a $\max\{\delta,\frac{6d\delta}{\lambda_{\min,1}}\}$-inclusion (we trivially have that $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is Hermitian preserving). However, as
@@
Basically, we simply apply Proposition 8 to $\widehat{\mathcal{F}}^{\otimes k}:\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes k})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes k})$} and $\widehat{\mathcal{G}}^{\otimes n}:\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})$, with $\widehat{\mathcal{F}}$ and $\widehat{\mathcal{G}}$ the reduced channels respectively associated to $\mathcal{F}$ and $\mathcal{G}$. Just note that the dimension of $\mathcal{H}_{\mathcal{F}}^{\otimes k}$ is $d^{k}$, with $d=\dim(\mathcal{H}_{\mathcal{F}})$. Thus, by Proposition 8, if the hypotheses of Theorem 11 are met, for all $p\in[1,+\infty]$, we have:
$\|\lambda(\mathcal{F}^{\otimes k})\|_{p}=\|\lambda(\widehat{\mathcal{F}}^{\otimes k})\|_{p}\leq\|\lambda(\widehat{\mathcal{G}}^{\otimes n})\|_{p}=\|\lambda(\mathcal{G}^{\otimes n})\|_{p}.$
This implies, by the compatibility of the $\ell_{p}$-norms of the shape vector with the tensor product, that